OpenStudy (anonymous):

I know this sounds stupid but it is hurting my brain. Ron made 32 foul shots out of 40 attempts. What is the probability that he will make his next shot?

OpenStudy (anonymous):

Assuming he continues with this success ratio; he has scored 8 shots out of 40 attempts, (40-32) hence the probability that he will score the next will be = 8/40 = 1/5 (after simplification)

OpenStudy (anonymous):

8/40 - Successful shots. So 8/40 = 1/5. The probability is 1/5 that he would make his next shot.

Directrix (directrix):

If by "What is the probability that he will make his next shot?" you mean the probability of making his next *foul shot*, empirical probability is, to my thinking, 32/40.= 4/5

OpenStudy (anonymous):

Thanks for all your help!